# Text Injection for Neural Contextual Biasing

### v1

@@ Abstract
Neural contextual biasing effectively improves automatic speech recognition (ASR) for crucial phrases within a speaker’s context, particularly those that are infrequent in the training data. This work proposes contextual text injection (CTI) to enhance contextual ASR. CTI leverages not only the paired speech-text data, but also a much larger corpus of unpaired text to optimize the ASR model and its biasing component. Unpaired text is converted into speech-like representations and used to guide the model’s attention towards relevant bias phrases. Moreover, we introduce a contextual text-injected (CTI) minimum word error rate (MWER) training, which minimizes the expected WER caused by contextual biasing when unpaired text is injected into the model. Experiments show that CTI with 100 billion text sentences can achieve up to 43.3% relative WER reduction from a strong neural biasing model. CTI-MWER provides a further relative improvement of 23.5%.
@@ 2.2 Neural Contextual Biasing with Deferred NAM
Upon removing red lines and red boxes, Fig. 1 dipicts a speech transducer [37, 26] with a neural biaser, deferred NAM [18]. The transducer consists of an acoustic encoder, a decoder and a joint network. The acoustic encoder transforms input speech features $\mathbf{X}$ into acoustic embeddings. To recognize named entities in speaker’s context, each speech utterance $\mathbf{X}$ is assigned $N$ bias phrases $Z=\{z_{1},...,z_{N}\}$. $z_{n}$ may or may not exist in $\mathbf{Y}$. Deferred NAM performs 2-pass hierarchical biasing at layer $l$ of the acoustic encoder. In Pass 1, we retrieve the K most relevant biasing phrases based on the correlations between acoustic embeddings $\mathbf{H}_{l}$ and bias phrase embeddings $\mathbf{P}$, followed by a max-pooling over $T$:
@@ 3 Contextual Text Injection
Figure 1: Contextual text injection for a speech transducer model. Blue and red lines show the forward propagation paths for paired audio-transcript data and unpaired text, respectively.
@@
The text-injected neural biasing model is trained to jointly minimize the two ASR losses and two retrieval losses computed from audio-transcript paired data $\mathcal{D}$ and the unpaired text $\mathcal{D}^{\text{text}}$:

### v2

@@ Abstract
Neural contextual biasing effectively improves automatic speech recognition (ASR) for crucial phrases within a speaker’s context, particularly those that are infrequent in the training data. This work proposes contextual text injection (CTI) to enhance contextual ASR. CTI leverages not only the paired speech-text data, but also a much larger corpus of unpaired text to optimize the ASR model and its neural biasing component. Unpaired text is converted into speech-like representations and used to guide the model’s attention towards relevant bias phrases. Moreover, we introduce a contextual text-injected (CTI) minimum word error rate (MWER) training, which minimizes the expected WER caused by contextual biasing when unpaired text is injected into the model. Experiments show that CTI with 100 billion text sentences can achieve up to 43.3% relative WER reduction from a strong neural biasing model. CTI-MWER provides a further relative improvement of 23.5%.
@@ 2.2 Neural Contextual Biasing with Deferred NAM
Upon removing red lines and red boxes, Fig. 1 depicts a speech transducer [37, 26] with a neural biaser, deferred NAM [18]. The transducer consists of an acoustic encoder, a decoder and a joint network. The acoustic encoder transforms input speech features $\mathbf{X}$ into acoustic embeddings. To recognize named entities in speaker’s context, each speech utterance $\mathbf{X}$ is assigned $N$ bias phrases $Z=\{z_{1},...,z_{N}\}$. $z_{n}$ may or may not exist in $\mathbf{Y}$. Deferred NAM performs 2-pass hierarchical biasing at layer $l$ of the acoustic encoder. In Pass 1, we retrieve the K most relevant biasing phrases based on the correlations between acoustic embeddings $\mathbf{H}_{l}$ and bias phrase embeddings $\mathbf{P}$, followed by a max-pooling over $T$:
@@ 3 Contextual Text Injection
Figure 1: Contextual text injection (CTI) for a speech transducer with neural biasing. Blue and red arrows show the forward propagation paths for paired audio-transcript data and unpaired text, respectively.
@@
The speech transducer, phoneme encoder, and neural biaser are jointly trained from scratch to minimize two ASR losses and two retrieval losses using audio-transcript paired data $\mathcal{D}$ and unpaired text $\mathcal{D}^{\text{text}}$: